Please provide a

1) Sales budget, including a schedule of expected cash collections.

2) Production budget.

Kimmel Gnomes, located in South Dakota, may be the only other manufacturer of high quality clay garden gnomes that are hand fFirst Third Fourth Quarter Ending FG inventory in Gnomes as a % of the next quarters budgeted sales 26% Ending FG inventoryVariable selling and administrative expenses are estimated to be $14.00 per gnome sold. Fixed selling and administrative expeLARGE GNOME DIVISION - KIMMEL GNOMES BUDGETED BALANCE SHEET JUNE 30, 2018 ASSETS LIABILITIES & EQUITY Accounts Cash $208,000
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Kimmel Gnomes Note
Sales Budget First Second Third Fourth Total
Sales (units)         44,000.00         82,000.00         42,000.00         76,000.00         244,000.00 A
Sell Price               107.00               107.00               107.00               107.00 B
Sales Budget 4,708,000.00 8,774,000.00 4,494,000.00 8,132,000.00 26,108,000.00 C=A*B
Collection Budget First Second Third Fourth Total
Sales Budget    4,708,000.00    8,774,000.00    4,494,000.00    8,132,000.00 26,108,000.00 See C
Collection of Accounts receivable of June 30    1,230,000.00      1,230,000.00 D= Taken from Balance Sheet
75% collected in same month    3,531,000.00    6,580,500.00    3,370,500.00    6,099,000.00 19,581,000.00 E= C*75%
25% collected in next month    1,177,000.00    2,193,500.00    1,123,500.00      4,494,000.00 F= 725% of C of previous month
Expected cash collections 4,761,000.00 7,757,500.00 5,564,000.00 7,222,500.00 25,305,000.00 G=D+E+F
Production Budget First Second Third Fourth Total
Sales (units)         44,000.00         82,000.00         42,000.00         76,000.00         244,000.00 See A
Add: Desired Ending Inventory         21,320.00         10,920.00         19,760.00         23,000.00 H= 26% of A of next month. For Fourth quarter 23,000 units is given in the sum.
Total units needed         65,320.00         92,920.00         61,760.00         99,000.00 I= A+H
Less: Expected Beginning Inventory                        -           21,320.00         10,920.00         19,760.00 J= 26% of A. For First quarter it is given the there is no finished goods inventory in the sum.
Production Budget         65,320.00         71,600.00         50,840.00         79,240.00         267,000.00 K=I-J